Project

General

Profile

NDOS Commissioning Meeting 2011/05/26

Attendance: Jon, Nick, Susan, Zeb, Jarek, Mark, Mark, Tricia, Gavin, Art, Minerba, Peter

Agenda:

  1. DataCheck Update
  2. NDOS Status and Channel Recovery Plans
  3. DAQ/Electronics Update & Plans
  4. DCS Update & Plans
  5. DataCheck Update
  6. AOB (All)

DataCheck Update

  • FEB-shutoff problem investigation continued. Correlated DAQ health monitoring data to FEB shutoff times.

NDOS Status

  • Jon:
    • uptime (run>30 minutes long and all 9 dcms in the readout) during the last week was ~67.4%
    • active channel fraction this past week:
      • whole-detector hovered around 70%
      • fiducial volume hovered around 77%
    • NuMI & BNB POT exposure plots are on the Commissioning Plots wiki . These are currently being updated by hand, but will soon be automated.
  • Mat:
    • 15 more cable-ectomized ribbons were installed on top of block 2. Summary is ~1/3 responded positively, including what was done previously. Mat will send around results later. Mark: after these 30, how many more? Mat: this is pretty much it.
  • Denis:
    • TECC cooling studies. Some FEBs still see significant noise (16% of all APDS). Tried to put more shielding and covering FEB cover with copper tape. Improvement seen on 3/5 that they worked on. Shielding reduced noise levels by ~3x on one of these.
    • Inspected some APDs that went bad after cooling on top of diblock 3. 5/6 were found to have o-ring out of groove (assembly was tight). Those 5 were removed from the detector. The 6th had a damaged o-ring, and the APD looked dirty, but no sign of condensation (dirt could be masking condensation marks).

DAQ/Electronics Update & Plans

  • Peter:
    • some improvements made as a result of stress tests. Increased buffer size in BNEVBs. Transferring data to BNEVBs has been taking longer than usual, and warning messages have been appearing in the message viewer (may have gone away now).

DCS Update & Plans

AOB (All)